# GLASS Flows: Transition Sampling for Alignment of Flow and Diffusion Models

### v1

@@ Appendix C Extended Related Work
Previous works have explored sampling from the posterior $p_{1|t}$ in other ways. For example, De Bortoli et al. (2025) explore learning the posterior posterior $p_{1|t}$ via a generative model trained via scoring rules allowing it to sample discrete-time transitions. In our experiments, we have also explored first sampling $z\sim p_{1|t}$ via GLASS Posterior Flows and then taking a conditional transition $p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t},z)$. However, we found it to lead to significantly worse performance than going “directly” to $x_{t^{\prime}}$ by sampling a GLASS transition (note that our setting is different as we reparameterize an existing model instead of training a new one as in (De Bortoli et al., 2025; Shaul et al., 2025)). Similarly, Gaussian mixture flow matching (Chen et al., 2025) approximates the posterior $p_{1|t}$ via a Gaussian mixture. This induces more stochasticity into a flow model and is shown to lead to performance improvements. Further, our method shares ideas from restart sampling Xu et al. (2023b), in that a variable is noised and then denoised via a flow again. However, our method is different in that the new noisy variable $\bar{x}_{s}$ still depends on $x_{t}$ up to $s=1$. Further, our method does not really go back in time (i.e. it does not restart, the time $t^{*}$ where the neural network is queried can monotonically increase). Finally, our method is not approximate but theoretically exact.
@@ D.2 Sampling Results
D.3 Reward Guidance
@@
D.4 Further Results for Feynman-Kac-Steering

### v2

@@ Appendix C Extended Related Work
Previous works have explored sampling from the posterior $p_{1|t}$, i.e. instead of just learning the mean via the denoiser actually sample from the posterior (De Bortoli et al., 2025; Elata et al., 2024; Chen et al., 2025a). For example, De Bortoli et al. (2025) explore learning the posterior posterior $p_{1|t}$ via a generative model trained via scoring rules allowing it to sample discrete-time transitions. In our experiments, we have also explored first sampling $z\sim p_{1|t}$ via GLASS Posterior Flows and then taking a conditional transition $p_{t^{\prime}|t}(x_{t^{\prime}}|x_{t},z)$. However, we found it to lead to significantly worse performance than going “directly” to $x_{t^{\prime}}$ by sampling a GLASS transition (note that our setting is different as we reparameterize an existing model instead of training a new one as in (De Bortoli et al., 2025; Shaul et al., 2025)). Similarly, Gaussian mixture flow matching (Chen et al., 2025a) approximates the posterior $p_{1|t}$ via a Gaussian mixture. This induces more stochasticity into a flow model and is shown to lead to performance improvements. Further, our method shares ideas from restart sampling Xu et al. (2023b), in that a variable is noised and then denoised via a flow again. However, our method is different in that the new noisy variable $\bar{x}_{s}$ still depends on $x_{t}$ up to $s=1$. Further, our method does not really go back in time (i.e. it does not restart, the time $t^{*}$ where the neural network is queried can monotonically increase). Finally, our method is not approximate but theoretically exact.
@@
Detailed discussion of Transition Matching (TM) (Shaul et al., 2025).
TM is a general pre-training framework for learning inner flow matching models. Due to changes in the neural network architecture and the focus on pre-training, TM and GLASS Flows are different, complementary methods. However, the theoretically optimal transitions are indeed closely related as we explain here in more detail. Specifically, we discuss here the DTM supervision process (Shaul et al., 2025). We write it here via a continuous time variable $0\leq t\leq 1$ using the convention of our work. In TM, the intermediate times are sampled via the CondOT probability path
$X_{t}=t\cdot X_{1}+(1-t)X_{0},\quad X_{1}=z\sim p_{\rm{data}},X_{0}\sim\mathcal{N}(0,I_{d})$
The training target is $Y=X_{1}-X_{0}$, i.e. TM learns a flow matching model to sampling from the conditional distribution of $Y|X_{t}=x_{t}$. Note that for fixed $X_{t}=x_{t}$, it holds that
$X_{0}=\frac{x_{t}-tX_{1}}{1-t}\quad\Rightarrow\quad Y=X_{1}-X_{0}=\frac{X_{1}-x_{t}}{1-t}$
Hence, sampling $Y$ or sampling $X_{1}$ is equivalent - one can transform each variable into one another. Further, let us define the probability path for the TM model conditioned on $x_{t}$:
$\tilde{X}_{s}=$ $\tilde{\alpha}_{s}Y+\tilde{\sigma}_{s}\epsilon\quad\epsilon\sim\mathcal{N}(0,I_{d})$
$=$ $\tilde{\alpha}_{s}\frac{X_{1}-x_{t}}{1-t}+\tilde{\sigma}_{s}\epsilon,\quad\epsilon\sim\mathcal{N}(0,I_{d})$
$=$ $\underbrace{\frac{\tilde{\alpha}_{s}}{1-t}X_{1}+\tilde{\sigma}_{s}\epsilon}_{=:\bar{X}_{s}}-\frac{\tilde{\alpha}_{s}}{1-t}x_{t}$
$=$ $\bar{X}_{s}-\frac{\tilde{\alpha}_{s}}{1-t}x_{t}$
where $\tilde{\alpha}_{s},\tilde{\sigma}_{s}$ are schedulers for the inner TM model, i.e. Shaul et al. (2025) choose $\tilde{\alpha}_{s}=s$ and $\tilde{\sigma}_{s}=1-s$. The variable $\bar{X}_{s}$ is distributed according to the GLASS probability path in eq. 18 if we chose $t^{\prime}=1$ and $\bar{\alpha}_{s}=\tilde{\alpha}_{s}/(1-t)$ and $\bar{\sigma}_{s}=\tilde{\sigma}_{s}$:
$p_{s}(\bar{x}_{s}|x_{t},z)=$ $\mathcal{N}(\bar{x}_{s};\bar{\alpha}_{s}z,\bar{\sigma}^{2}_{s}I_{d})$
where we used that $\bar{\gamma}=0$ in this case. Therefore, if we simulate a GLASS trajectory with these parameters ($t^{\prime}=1,\bar{\sigma}_{s}=\tilde{\sigma}_{s},\bar{\alpha}_{s}=\tilde{\alpha}_{s}/(1-t)$), we obtain that the transformed GLASS trajectory
$\tilde{X}_{s}=\bar{X}_{s}-\frac{\tilde{\alpha}_{s}}{1-t}x_{t}$
is a trajectory obtained from TM/DTM. This elucidates the connection between DTM TM and GLASS Flows.
Detailed discussion of TADA (Chen et al., 2025b).
We provide an extended discussion of how the TADA method relates to this work. Specifically, we rewrite a simplified argument from (Chen et al., 2025b) in the notation of this work to showcase the connection. Specifically, we focus here on a simple case of $N=2$, i.e. augmenting the data space $z\in\mathbb{R}^{d}$ with a single variable $p\in\mathbb{R}^{d}$, and where the terminal point $\tilde{\mathbf{x}}_{1}=(z,p_{1})$ in the state space is sampled independently, i.e. $z\sim p_{\rm{data}}$ and $p\sim\mathcal{N}(0,I_{d})$ (for the argument in full generality, we refer to (Chen et al., 2025b)). The probability path in this case is given by
$\tilde{\mathbf{x}}_{t}=(\mu_{t}\otimes I_{d})\tilde{\mathbf{x}}_{1}+(L_{t}\otimes I_{d})\mathbf{\epsilon},\quad\mathbf{\epsilon}=(\epsilon_{1},\epsilon_{2})^{T},\epsilon_{1},\epsilon_{2}\sim\mathcal{N}(0,I_{d})$
The core realization connecting it to our work is now that that $\tilde{\mathbf{x}}_{t}=(x_{t},p_{t})$ consists of two components $x_{t}$ and $p_{t}$ that are both in themselves noisy (correlated) Gaussian measurements of $z$. This follows simply from the fact $p_{1}$ is Gaussian and $\epsilon=(\epsilon_{1},\epsilon_{2})$ is Gaussian and linear transformations of Gaussian are again Gaussian. Hence, we are in a similar setting as in section 4.2.1. In fact, Chen et al. (2025b) applied a similar argument to recover the denoiser from a pre-trained flow matching or diffusion model as in section 4.2.1 (see (Chen et al., 2025b, Proposition 3.1)). This showcases the close connection of the mathematical principles enabling (Chen et al., 2025b) and this work.
@@ D.2 Sampling Results
Figure 6: Evaluating diversity of samples from various sampling scheme. All 3 sampling approachess (DDPM, ODE, GLASS) show very similar results for diversity. This is consistent with theory as all 3 approaches should sample from the same distribution. We evaluate the model with 100 total NFEs on GenEval prompts (we take more NFEs than in fig. 3 to reduce discretization error of the SDE). We take 8 samples per prompt and measure the average DreamSim (Fu et al., 2023) similarity between two samples (error bars equal average standard deviation of samples of prompt).
D.3 Ablation over correlation parameter $\rho$
We perform further experiments in this section ablating the correlation parameter $\rho$. We choose two strategies for ablation: First, a constant correlation parameter $\rho$ is chosen across all transitions from $t\to t^{\prime}$ independent of $t,t^{\prime}$. Second, we choose $\rho$ as time-varying based on the DDPM schedule (see proposition 1):
$\rho=\left(\frac{\alpha_{t}\sigma_{t^{\prime}}}{\sigma_{t}\alpha_{t^{\prime}}}\right)^{\kappa}$
and we ablate over the parameter $\kappa\geq 0$. We use the FLUX model and measure GenEval performance for prompt adherence/generation quality and DreamSim diversity (Fu et al., 2023) as a measure of diversity. We present results in fig. 7. The most striking results is that GLASS Flows achieves ODE-level performance for almost all correlation schedules and differences of performance are relatively minor. A constant correlation schedule of $\rho=0.4$ performs best in GenEval performance and has relatively high sample diversity. Therefore, we choose $\rho=0.4$ in subsequent experiments. For the ablation over $\kappa$, DDPM ($\kappa=1$) performs very high (just diversity is slightly higher for $\kappa=2$). Due to the wide use of DDPM and its theoretical importance as a time-reversal, we also use DDPM in subsequent experiments.
Figure 7: Ablation experiment of correlation schedule $\rho$. Left: Constant correlation $\rho$ across all transitions. Right: Time-dependent correlation schedule given by $\rho=\left(\frac{\alpha_{t}\sigma_{t^{\prime}}}{\sigma_{t}\alpha_{t^{\prime}}}\right)^{\kappa}$ - note that $\kappa=1$ corresponds to the DDPM schedule (see proposition 1).
D.4 Reward Guidance
@@
D.5 Further Results for Feynman-Kac-Steering